The Sesnie Company is preparing its cash payments budget. The following items relate to cash payments the company anticipatesa. The company pays for 50% of its direct materials purchases in the month of purchase and the remainder the following month.
Answer #1
Budgeted Cash Payments
April May June Quarter
Current Month Purchases (50%) $    67,000 $       64,000 $          73,000
Previous Month Purchases (50%) $    55,000 $       67,000 $          64,000
Total Cash Payments for Purchases $ 122,000 $    131,000 $        137,000 $       390,000
Direct Labor $    56,000 $       66,000 $          81,000 $       203,000
Manufacturing Overhead $    53,600 $       69,600 $          93,600 $       216,800
Operating Expenses $    59,200 $       76,200 $          73,200 $       208,600
New Server $         2,000 $            2,000
Tax Payment $          12,500 $         12,500
Total Cash Payments $ 290,800 $    344,800 $        397,300 $   1,032,900

Working

Budgeted Cash Payments
April May June
Current Month Purchases (50%) =134000*50% =128000*50% =146000*50%
Previous Month Purchases (50%) =110000*50% =134000*50% =128000*50%
Total Cash Payments for Purchases
Direct Labor 56000 66000 81000
Manufacturing Overhead =56000*160%-36000 =66000*160%-36000 =81000*160%-36000
Operating Expenses =70000-8000-2800 =87000-8000-2800 =84000-8000-2800
New Server 2000
Tax Payment 12500